OpenStudy (anonymous):

If g(x) = 3(x - 2), find the value of x if g(x) = 6. 12 2.6 8 4

OpenStudy (anonymous):

@zBrandz23

OpenStudy (jdoe0001):

g(x) = 3(x - 2), g(x)=6 thus 6 = 3(x-2) <---- solve for "x"

OpenStudy (anonymous):

4=3x 1.1 =x ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@zBrandz23

jimthompson5910 (jim_thompson5910):

you can't go from 6 = 3(x-2) to 4 = 3x because a) the -2 is not on the side with the 6 b) the -2 is buried in the parenthesis

OpenStudy (anonymous):

So what do i do then?

jimthompson5910 (jim_thompson5910):

you first need to distribute 6 = 3(x-2) 6 = 3x-6

jimthompson5910 (jim_thompson5910):

whats next?

OpenStudy (anonymous):

move -6 to the other side making it +6 so 12=3x so 4=x

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

:D thanks soooo much :D can i ask another on another question i'll tag you in

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure, one more